OpenStudy (anonymous):

7a squared+8ab 4p-pq-5pr 6st+5su+7sv 4pq squared+5qr+3pqr 9aS quared b+ab squared-8ab 7t squared u squared+3tu sqaured v+8tuw 20abc squared-8a squared bc +12ab squared c 30pqrs +20q squared rs- 70 qrs

OpenStudy (anonymous):

write it with equation forum

OpenStudy (anonymous):

how do yuo do the little 2?

OpenStudy (anonymous):

the little 2 ? didn't understand

OpenStudy (anonymous):

the squared little 2

OpenStudy (aroub):

just do like that, ^2

OpenStudy (anonymous):

\[\sqrt{2}\] or \[[2^2]\]

OpenStudy (anonymous):

help me man

jhonyy9 (jhonyy9):

wath do you like ?

jhonyy9 (jhonyy9):

\[7a ^{2}+8ab\]

jhonyy9 (jhonyy9):

=a(7a+8b)

jhonyy9 (jhonyy9):

4p-pq-5pr=p(4-q-5r)

jhonyy9 (jhonyy9):

s(6t+5u+7v)